# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 1 Introduction
The mathematical theory of fair item allocation among multiple agents has practical applications in scenarios like inheritance and partnership dissolutions. When agents have equal entitlements, as in inheritance cases, each agent naturally expects their allocated portion to be at least as good as others’. An allocation satisfying this requirement is called envy-free (EF).
When the items available for allocation are all indivisible, an envy-free allocation might not exist. A solution often applied in practice is to use money to compensate for the envy.
@@
It is common to assume that the agents are quasilinear. This means that their utility equals their total value for the items they receive, plus the amount of money they receive (which may be positive or negative). Alternatively, one can consider a hypothetical third-party, who is willing to subsidize the process such that all agents receive a non-negative amount, and ask what is the minimum amount of subsidy that is required to attain envy-freeness.
@@
The subsidy minimization problem was first studied by Halpern and Shah [15]. They showed that for any given allocation, there exists a permutation of the bundles that can be made envy-free with subsidies. The total required subsidy is at most $(n-1)mV$, where $m$ is the number of items, $n$ the number of agents, and $V$ the maximum item value for an agent, and this bound is tight in the worst case. Brustle et al. [7] considered the case in which the allocation is not given, but can be chosen. They presented an algorithm that finds an envy-freeable allocation through iterative maximum matching, requiring a total subsidy of at most $(n-1)V$, which is tight too.
@@
This paper extends previous work on subsidy minimization by considering agents with different entitlements. In partnership dissolutions, partners often hold varying numbers of shares, entitling them to different proportions of the asset. Each agent expects to receive at least the same ”value per share” as others.
For example, if agent $i$ has twice the entitlement of agent $j$, $i$ expects a bundle worth at least twice as much as $j$’s.
@@
Formally, an allocation is called weighted envy-free (WEF) (see e.g. [20, 32, 18]) if for every two agents $i$ and $j$, $\frac{1}{w_{i}}$ times the utility that $i$ assigns to his own bundle is at least as high as $\frac{1}{w_{j}}$ the utility that $i$ assigns to the bundle of $j$, where $w_{i}$ is $i$’s entitlement and $w_{j}$ is $j$’s entitlement. To illustrate the difficulty in this generalized setting, we show that both results mentioned above [15, 7] fail when agents have different entitlements.
Example 1.1.
@@
We will show that, contrary to the result of Halpern and Shah [15], there exists a division of items where no permutation satisfies WEF. Consider the bundles $A_{1}=\{o_{1}\}$ and $A_{2}=\{o_{2}\}$, where $i_{1}$ receives $A_{1}$ and $i_{2}$ receives $A_{2}$. Let $s_{1}$ and $s_{2}$ represent the subsidies for $i_{1}$ and $i_{2}$, respectively. The utility of $i_{1}$ for their own bundle is $5+s_{1}$, and for $i_{2}$’s bundle, it is $8+s_{2}$. To satisfy WEF, we need: $\frac{5+s_{1}}{1}\geq\frac{7+s_{2}}{10}$, which implies $s_{2}\leq 43+10s_{1}$. Similarly, for agent $i_{2}$, WEF requires: $\frac{8+s_{2}}{10}\geq\frac{10+s_{1}}{1}$, which implies $s_{2}\geq 92+10s_{1}$. These two conditions are contradictory, so no subsidies can make this allocation WEF. Next, consider the permutation where $i_{1}$ receives $A_{2}$ and $i_{2}$ receives $A_{1}$. In this case, WEF requires: $\frac{7+s_{1}}{1}\geq\frac{5+s_{2}}{10}$, which implies $s_{2}\leq 65+10s_{1}$. and for $i_{2}$: $\frac{10+s_{2}}{10}\geq\frac{8+s_{1}}{1}$, which implies $s_{2}\geq 70+10s_{1}$. Again, these conditions are contradictory, proving that no permutation of bundles satisfies WEF.
@@ 1.1 Related Work
Steinhaus [22] initiated fair allocation with the cake-cutting problem, followed by Foley’s [13] advocacy for envy-free resource allocation. Challenges with indivisible items were outlined by Schmeidler and Yaari [21].
@@
Particularly relevant to our work is a recent work by Viswanathan and Zick [25], who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms.
@@
The concept of compensating an indivisible resource allocation with money has been explored in the literature ever since Demange, Gale, and Sotomayor [12] introduced an ascending auction for envy-free allocation using monetary payments for unit demand agents.
In addition to their results for additive valuations mentioned earlier, Brustle et al. [7] also study the more general class of monotone valuations. They demonstrate that a total subsidy of $2(n-1)^{2}V$ is sufficient to guarantee the existence of an envy-freeable allocation. Kawase el al. [16] improved this bound to $\frac{n^{2}-n-1}{2}$.
Caragiannis and Ioannidis [8] developed an algorithm that approximates the minimum subsidies with any required accuracy for a constant number of agents, though with increased running time. However, for a super-constant number of agents, they showed that minimizing subsidies for envy-freeness is both hard to compute exactly and difficult to approximate.
Aziz [2] presented a sufficient condition and an algorithm to achieve envy-freeness and equitability (every agent should get the same utility) when monetary transfers are allowed for agents with quasi-linear utilities and superadditive valuations (positive or negative).
Barman et al. [6] studied agents with dichotomous valuations (agents whose marginal value for any good is either zero or one), without any additivity requirement. They proved that, for $n$ agents, there exists an allocation that achieves envy-freeness with total required subsidy of at most $n-1$, which is tight even for additive valuations.
Goko et al. [14] study an algorithm for an envy-free allocation with subsidy, that is also truthful, when agents have submodular binary valuations. The subsidy per agent is at most $V=1$. Their algorithm works only for agents with equal entitlements.
@@
In the past few years, several researchers have examined a more general model in which different agents may have different entitlements, included weighted fairness models like weighted maximin share fairness (WMMS) and weighted proportionality up to one item (WPROP1) [9, 5, 3]. Chakraborty et al. [18] established maximum weighted Nash welfare (MWNW) satisfies Pareto optimality and introduced a weighted extension of EF1. Suksompong and Teh [23] demonstrated MWNW properties under binary additive valuations and its polynomial-time computability. They further extended these findings to various valuation types.
@@
Wu et al. ([31]) presented a polynomial-time algorithm for computing a PROP allocation of chores among agents with additive valuations, with total subsidy at most $\frac{nV}{4}$, which is tight. For agents with different entitlements, they compute a WPROP allocation with total subsidy at most $\frac{(n-1)V}{2}$. In a subsequent work ([30]), they further improved this bound to $(\frac{n}{3}-\frac{1}{6})V$.
@@
Recent studies have considered matroid-rank valuation (binary submodular). Montanari et al. [19] introduce a new family of weighted envy-freeness notions based on the concept of transferability and provide an algorithm for computing transferable allocations that maximize welfare. Babaioff et al. [4] design truthful allocation mechanisms that maximize welfare and are fair. Particularly relevant to our work is a recent work by Viswanathan and Zick [25], who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms. Later, Viswanathan and Zick in [26] generalize the Yankee Swap algorithm to efficiently compute allocations that maximize any fairness objective, called General Yankee Swap. We show that, under binary additive valuations, the general Yankee Swap algorithm can be adapted to agents with different entitlements. The adaptation is shown in Section 5.
@@ 1.2 Our Results
In Section 3.1, we assume that an allocation is given. As shown in Example 1.1, there are instances in which no rearrangement of bundles yields a WEF-able allocation. We prove a necessary and sufficient condition under which the allocation is WEF-able. We show that, when the allocation is WEF-able, a total subsidy of $(\frac{W}{w_{1}}-1)mV$ is sufficient to make it WEF, and prove that this bound is tight in the worst case.
This raises the question of whether a weighted-envy-free allocation with subsidy always exists? We provide an affirmative answer in Section 3.2. We show how to compute for agents with integer weights, a WEF-able allocation with total subsidy at most $\left(W-w_{1}\right)$.
Our algorithm extends Brustle et al. [7] algorithm, which in the unweighted setting attains total subsidy at most $(n-1)V$. With equal entitlements, we can normalize the weights to 1, maintaining the same bounds as in the unweighted case.
Following [15], in addition to the setting of general additive valuations, we study the setting in which agents have identical additive valuations (Section 4), and the setting in which agents have binary additive valuations (Section 5).
For identical additive valuations, we demonstrate an efficient computation of a WEF allocation with subsidies capped at $V$ for each agent, and a total subsidy bounded by $(n-1)V$, which is identical to the unweighted setting. We also prove that this bound is tight in the worst case.
@@
In Section 5, under binary-additive valuations, we show how to modify the General Yankee Swap algorithm of [26] to efficiently compute a WEF-able allocation with subsidies not exceeding $\frac{w_{i}}{w_{1}}$ per agent $i\in N$, and a total subsidy bounded by $\frac{W-w_{1}}{w_{1}}=\frac{W}{w_{1}}-1$, if all agents’ weights are integer. This bound reduces to an identical overall subsidy of $n-1$ in the unweighted setting.
@@
Character- ization of WEF Allocation (1) No positive cost cycles, (2) USW maximi- zation [15] No positive-cost cycles. (3.2)
Permut- ation of a given allocation, that maximizes sum of values Always EF-able [15] Not necessarily WEF-able (3.3) Always WEF-able. (4.3) For non- redundant allocation: Always WEF-able (5.5)
Total subsidy upper bound $(n-1)V$ [15] $(W-w_{1})V$ (3.17) $(n-1)V$ (4.7) $\frac{W}{w_{1}}-1$ (5.8)
Subsidy bound for a given allocation $(n-1)mV$ [15] $\left(\frac{W}{w_{1}}-1\right)mV$ (3.5)
Total subsidy lower bound $(n-1)V$ [15] $\left(\frac{W}{w_{1}}-1\right)V$ (3.9) $(n-1)V$ (4.1) $\frac{W}{w_{2}}-1$ (5.1)
Remark 1.2.
@@ 2 Preliminaries
We assume that agents are quasilinear, so that the utility of each agent is $u_{i}(A_{i},s_{i}):=v_{i}(A_{i})+s_{i}$.
We assume that the valuation functions $v_{i}$ are normalized, that is $v_{i}(\emptyset)=0$ for all $i\in N$ (Suksompong and Teh. [24]); and additive, that is $v_{i}(S)=\sum_{o\in S}v_{i}(o)$ for all $S\subseteq M,i\in N$. In some sections of this paper, we make some additional assumptions:
@@
$\frac{u_{i}(A_{j})}{w_{j}}-\frac{u_{i}(A_{i})}{w_{i}}=\frac{v_{i}(A_{j})+s_{j}}{w_{j}}-\frac{u_{i}(A_{i})+s_{i}}{w_{i}}$
@@
In the setting without money, WEF can be relaxed by allowing envy up to an upper bound that depends on the item values. We employ the generalization of allowable envy proposed by Chakraborty et al. [10]:
Definition 2.2.
(Chakraborty et al.[10]). For $x,y\in[0,1]$, an allocation $A$ is said to satisfy WEF(x,y) if for any $i,j\in N$, there exists $B\subseteq A_{j}$ with $|B|\leq 1$ such that $\frac{v_{i}(A_{i})+yv_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j})-xv_{i}(B)}{w_{j}}$, where $x+y=1$.
@@
$cost_{A}(i,j):=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}$
@@
The previous work of Halpern and Shah [15] provides sufficient and necessary conditions for an EF-able allocation in the unweighted setup. The following theorem considers a pre-determined allocation $A=(A_{1},\ldots,A_{n})$, where we aim to assign one bundle to each agent.
Theorem 2.3.
(Halpern and Shah [15]). In a setting with equal entitlements, the following statements are equivalent:
@@
    The allocation $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents, that is, for every permutation $\sigma$ of $[n]$ (a bijection $\sum:[n]\rightarrow[n]$), $\sum_{i\in N}v_{i}(A_{i})\geq\sum_{i\in N}v_{i}(A_{\sigma{i}})$.
@@
Halpern and Shah [15] have outlined a method to find a subsidy vector $\mathbf{s}$, ensuring envy-freeness while minimizing the required subsidy. Each agent’s subsidy is determined by the maximum cost path from that agent in the envy graph. This cost can be computed within strongly polynomial time.
@@
    When the allocation is given: the minimum subsidy required is $(n-1)mV$ in the worst case. [15].
@@
    When the allocation can be chosen: the minimum subsidy required is at least $(n-1)V$ in the worst case, even in the special case of binary valuations [15].

### v4

@@ 1 Introduction
The mathematical theory of fair item allocation among multiple agents has practical applications in scenarios like inheritance and partnership dissolutions. When agents have equal entitlements, as in inheritance cases, each agent naturally expects their allotment to be at least as good as others’. An allocation satisfying this requirement is called envy-free (EF).
@@
When the items available for allocation are indivisible, an EF allocation might not exist. A solution often applied in practice is to use money to compensate for the envy. In the recent literature, it is common to assume that a hypothetical third-party is willing to subsidize the process such that all agents receive a non-negative amount, and ask what is the minimum amount of subsidy required to attain envy-freeness.
It is common to assume that the agents are quasilinear. This means that their utility equals their total value for the items they receive, plus the amount of money they receive (which may be positive or negative).
@@
The subsidy minimization problem was first studied by Halpern and Shah [18]. They showed that, for any given allocation, there exists a permutation of the bundles that is envy-freeable (EF-able), that is, can be made envy-free with subsidies. The total required subsidy is at most $(n-1)mV$, where $m$ is the number of items, $n$ the number of agents, and $V$ the maximum item value for an agent, and this bound is tight in the worst case. Brustle et al. [7] considered the case in which the allocation is not given, but can be chosen. They presented an algorithm that finds an envy-freeable allocation through iterative maximum matching, requiring a total subsidy of at most $(n-1)V$, which is tight too.
@@
This paper extends previous work by considering agents with different entitlements, which we call weights. This extension is useful in partnership dissolutions, where partners often hold varying numbers of shares, entitling them to different proportions of the asset. In such cases, each agent expects to receive at least the same ”value per share” as others. For example, if agent $i$ has twice the entitlement of agent $j$, $i$ expects a bundle worth at least twice as much as $j$’s.
@@
Formally, an allocation is called weighted envy-free (WEF) (see e.g. Robertson and Webb [24], Zeng [36], Chakraborty et al. [9]) if for every two agents $i$ and $j$, $\frac{1}{w_{i}}$ times the utility that $i$ assigns to his own bundle is at least as high as $\frac{1}{w_{j}}$ times the utility that $i$ assigns to the bundle of $j$, where $w_{i}$ is $i$’s entitlement and $w_{j}$ is $j$’s entitlement.
Now, we define weighted envy-freeability (WEF-ablity), the key concept we propose, analogously to the unweighted case: an allocation is WEF-able if it can be made WEF with subsides. More presicely, an allocation is WEF-able if for every two agents $i$ and $j$, $\frac{1}{w_{i}}$ times the sum of the utility that $i$ assigns to his own bundle and the subsidy he receives is at least as high as $\frac{1}{w_{j}}$ times the sum of the utility that $i$ assigns to the bundle of $j$ and the subsidy $j$ receives, where $w_{i}$ (resp., $w_{j}$) is $i$’s (resp., $j$’s) entitlement.
Here, we assume quasi-linear utilities. To illustrate the difficulty in this generalized setting, we show that the results from Halpern and Shah [18], Brustle et al. [7] fail when agents have different entitlements.
[…]
@@
We will show that, contrary to the result of Halpern and Shah [18], there exists a division of items where no permutation satisfies WEF. Consider the bundles $A_{1}=\{o_{1}\}$ and $A_{2}=\{o_{2}\}$, where $i_{1}$ receives $A_{1}$ and $i_{2}$ receives $A_{2}$. Let $s_{1}$ and $s_{2}$ represent the subsidies for $i_{1}$ and $i_{2}$, respectively. The utility of $i_{1}$ for their own bundle is $5+s_{1}$, and for $i_{2}$’s bundle, it is $7+s_{2}$. To satisfy WEF, we need: $\frac{5+s_{1}}{1}\geq\frac{7+s_{2}}{10}$, which implies $s_{2}\leq 43+10s_{1}$. Similarly, for agent $i_{2}$, WEF requires: $\frac{8+s_{2}}{10}\geq\frac{10+s_{1}}{1}$, which implies $s_{2}\geq 92+10s_{1}$. These two conditions are contradictory, so no subsidies can make this allocation WEF. Next, consider the permutation where $i_{1}$ receives $A_{2}$ and $i_{2}$ receives $A_{1}$. In this case, WEF requires: $\frac{7+s_{1}}{1}\geq\frac{5+s_{2}}{10}$, which implies $s_{2}\leq 65+10s_{1}$, and for $i_{2}$: $\frac{10+s_{2}}{10}\geq\frac{8+s_{1}}{1}$, which implies $s_{2}\geq 70+10s_{1}$. Again, these conditions are contradictory, proving that no permutation of bundles satisfies WEF.
@@
Of course, since the unweighted case is equivalent to the weighted case where each weight $w_{i}=1/n$, all negative results from the unweighted setting extend to the weighted case. In particular, it is NP-hard to compute the minimum subsidy required to achieve (weighted) envy-freeness, even in the binary additive case (as shown in [halpern2019fair, Corollary 1]). Thus, following previous work, we develop polynomial-time algorithms that, while not necessarily optimal, guarantee an upper bound on the total subsidy.
@@ 1.1 Related Work
steinhaus1948problem initiated fair allocation with the cake-cutting problem, followed by foley1966resource advocacy for envy-free resource allocation. Challenges with indivisible items were outlined by schmeidler1971fair.
@@
Particularly relevant to our work is a recent work by viswanathan2022yankee, who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms.
@@
The concept of compensating an indivisible resource allocation with money has been explored in the literature ever since demange1986multi introduced an ascending auction for envy-free allocation using monetary payments for unit demand agents.
In addition to their results for additive valuations mentioned earlier, brustle2020one also study the more general class of monotone valuations. They demonstrate that a total subsidy of $2(n-1)^{2}V$ is sufficient to guarantee the existence of an envy-freeable allocation. kawase2024towards improved this bound to $\frac{n^{2}-n-1}{2}$.
caragiannis2021computing developed an algorithm that approximates the minimum subsidies with any required accuracy for a constant number of agents, though with increased running time. However, for a super-constant number of agents, they showed that minimizing subsidies for envy-freeness is both hard to compute exactly and difficult to approximate.
aziz2021achieving presented a sufficient condition and an algorithm to achieve envy-freeness and equitability (every agent should get the same utility) when monetary transfers are allowed for agents with quasi-linear utilities and superadditive valuations (positive or negative).
barman2022achieving studied agents with dichotomous valuations (agents whose marginal value for any good is either zero or one), without any additivity requirement. They proved that, for $n$ agents, there exists an allocation that achieves envy-freeness with total required subsidy of at most $n-1$, which is tight even for additive valuations.
goko2024fair study an algorithm for an envy-free allocation with subsidy, that is also truthful, when agents have submodular binary valuations. The subsidy per agent is at most $V=1$. Their algorithm works only for agents with equal entitlements.
@@
In the past few years, several researchers have examined a more general model in which different agents may have different entitlements, included weighted fairness models like weighted maximin share fairness (WMMS) and weighted proportionality up to one item (WPROP1) (chakraborty2021picking, babaioff2023fair, aziz2019weighted). mithun2021weighted established maximum weighted Nash welfare (MWNW) satisfies Pareto optimality and introduced a weighted extension of EF1. suksompong2022maximum demonstrated MWNW properties under binary additive valuations and its polynomial-time computability. They further extended these findings to various valuation types.
@@
wu2023one presented a polynomial-time algorithm for computing a PROP allocation of chores among agents with additive valuations, with total subsidy at most $\frac{nV}{4}$, which is tight. For agents with different entitlements, they compute a WPROP allocation with total subsidy at most $\frac{(n-1)V}{2}$. In a subsequent work (wu2024tree), they further improved this bound to $(\frac{n}{3}-\frac{1}{6})V$.
@@
Recent studies have considered matroid-rank valuation (binary submodular). montanari2024weighted introduce a new family of weighted envy-freeness notions based on the concept of transferability and provide an algorithm for computing transferable allocations that maximize welfare. babaioff2021fair design truthful allocation mechanisms that maximize welfare and are fair. Particularly relevant to our work is a recent work by viswanathan2022yankee, who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms. Later, viswanathan2023general generalize the Yankee Swap algorithm to efficiently compute allocations that maximize any fairness objective, called General Yankee Swap. We show that, under binary additive valuations, the general Yankee Swap algorithm can be adapted to agents with different entitlements. The adaptation is shown in Section 5.
@@ 1.2 Our Results
In Section 3.1, we assume that an allocation is given. As shown in Example 1, there are instances in which no rearrangement of bundles yields a WEF-able allocation. We prove a necessary and sufficient condition under which the allocation is WEF-able. We show that, when the allocation is WEF-able, a total subsidy of $(\frac{W}{w_{1}}-1)mV$ is sufficient to make it WEF, and prove that this bound is tight in the worst case.
This raises the question of whether a weighted-envy-free allocation with subsidy always exists? We provide an affirmative answer in Section 3.2. For agents with additive valuations, assuming all weights are integers, we show an upper bound that is independent of $m$: it is $\frac{W-w_{1}}{\operatorname{gcd}(\mathbf{w})}V$, where $\operatorname{gcd}(\mathbf{w})$ is the greatest common divisor of all weights — largest number $d$ such that $w_{i}/d$ is an integer for all $i\in N$
Our algorithm extends brustle2020one algorithm, which in the unweighted setting attains total subsidy at most $(n-1)V$. With equal entitlements, we can normalize the weights to 1, maintaining the same bounds as in the unweighted case.
Following halpern2019fair, in addition to the setting of general additive valuations, we study the setting in which agents have identical additive valuations (Section 4), and the setting in which agents have binary additive valuations (Section 5).
For identical additive valuations, we compute a WEF-able and WEF$(0,1)$ allocation with total subsidy at most $(n-1)V$, which is tight even in the unweighted case.
@@
In Section 5, under binary-additive valuations, we adapt the General Yankee Swap algorithm [viswanathan2023general] to compute a WEF-able and WEF$(0,1)$ allocation with total subsidy at most $\frac{W}{w_{1}}-1$, reducing to $n-1$ for equal weight.
@@
Character- ization of WEF Allocation (1) No positive cost cycles, (2) USW maximi- zation [halpern2019fair] No positive-cost cycles. (3.1)
Permut- ation of a given allocation, that maximizes sum of values Always EF-able [halpern2019fair] Not necessarily WEF-able (3.1) Always WEF-able. (4) For non- redundant allocation: Always WEF-able (5)
Total subsidy upper bound $(n-1)V$ [halpern2019fair] $\frac{W-w_{1}}{\operatorname{gcd}(\mathbf{w})}V$ (3.2) $(n-1)V$ (4) $\frac{W}{w_{1}}-1$ (5)
Subsidy bound for a given allocation $(n-1)mV$ [halpern2019fair] $\left(\frac{W}{w_{1}}-1\right)mV$ (3.1)
Total subsidy lower bound $(n-1)V$ [halpern2019fair] $\left(\frac{W}{w_{1}}-1\right)V$ (3.2) $(n-1)V$ (4) $\frac{W}{w_{2}}-1$ (5)
[…]
@@ 2 Preliminaries
We assume that agents are quasilinear, so that the utility of each agent is
$u_{i}(A_{i},s_{i}):=v_{i}(A_{i})+s_{i}.$
We assume that the valuation functions $v_{i}$ are normalized, that is $v_{i}(\emptyset)=0$ for all $i\in N$ (suksompong2023weighted); and additive, that is $v_{i}(S)=\sum_{o\in S}v_{i}(o)$ for all $S\subseteq M,i\in N$. In some sections of this paper, we make some additional assumptions:
@@
$\frac{u_{i}(A_{j})}{w_{j}}-\frac{u_{i}(A_{i})}{w_{i}}=\frac{v_{i}(A_{j})+s_{j}}{w_{j}}-\frac{u_{i}(A_{i})+s_{i}}{w_{i}}.$
@@
In the setting without money, WEF can be relaxed by allowing envy up to an upper bound that depends on the item values. We employ the generalization of allowable envy proposed by chakraborty2022weighted:
[…]
(chakraborty2022weighted). For $x,y\in[0,1]$, an allocation $A$ is said to satisfy WEF(x,y) if for any $i,j\in N$, there exists $B\subseteq A_{j}$ with $|B|\leq 1$ such that
$\frac{v_{i}(A_{i})+yv_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j})-xv_{i}(B)}{w_{j}}.$
@@
$cost_{A}(i,j):=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}.$
@@
The previous work of halpern2019fair provides sufficient and necessary conditions for an EF-able allocation in the unweighted setup. The following theorem considers a pre-determined allocation $A=(A_{1},\ldots,A_{n})$, where we aim to assign one bundle to each agent.
{theorem}
(halpern2019fair). In a setting with equal entitlements, the following statements are equivalent:
@@
    The allocation $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents, that is, for every permutation $\sigma$ of $[n]$ (a bijection $\sigma:[n]\rightarrow[n]$), $\sum_{i\in N}v_{i}(A_{i})\geq\sum_{i\in N}v_{i}(A_{\sigma{i}})$.
@@
halpern2019fair have outlined a method to find a subsidy vector $\mathbf{s}$, ensuring envy-freeness while minimizing the required subsidy. Each agent’s subsidy is determined by the maximum cost path from that agent in the envy graph. This cost can be computed within strongly polynomial time.
@@
    When the allocation is given: the minimum subsidy required is $(n-1)mV$ in the worst case [halpern2019fair].
@@
    When the allocation can be chosen: the minimum subsidy required is at least $(n-1)V$ in the worst case, even in the special case of binary valuations [halpern2019fair].